\begin{thebibliography}{99}


\bibitem{CDY05} C. Carlet, C. Ding, and J. Yuan, Linear codes from highly nonlinear functions and their
secret sharing schemes, IEEE Trans. Inform. Theory 51 (6) (2005) 2089--2102.

\bibitem{Delsarte} P. Delsarte, On subfield subcodes of modified Reed-Solomon codes,
IEEE Trans. Inform. Theory  21 (5) (1975) 575--576.

\bibitem{Ding12} C. Ding and J. Yang, Hamming weights in irreducible cyclic codes,
Discrete Mathematics 313 (4) (2013) 434--446.

\bibitem{Ding13} C. Ding and S. Ling, A $q$-polynomial approach to cyclic codes, Finite Fields Appl. 20 (3) (2013) 1--14.

\bibitem{DingSAM}C. Ding, Cyclic codes from some monomials and trinomials, SIAM J. Discrete Mathematics
27 (4) (2013) 1977--1994.

\bibitem{Ding-Heelseth} C. Ding and T. Helleseth, Optimal ternary cyclic codes from monomials, IEEE Trans. Inform. Theory 59 (9) (2013) 5898--5904.

\bibitem{DW05} C. Ding and X. Wang, A coding theory construction of new systematic authentication codes, {Theor. Comput. Sci.} 330 (1) (2005) 81--99.

\bibitem{DYT} C. Ding, Y. Yang, and X. Tang, Optimal sets of frequency hopping sequences from linear cyclic codes, {IEEE Trans. Inform. Theory} 56 (7) (2010) 3605--3612.

\bibitem{Dobb} H. Dobbertin, T. Helleseth, P. V. Kumar, and H. Martinsen, Ternary m-sequences with three-valued cross-correlation function: New decimations of Welch and Niho type, IEEE Trans. Inform.
Theory 47 (4) (2001) 1473--1481.

\bibitem{FLZ} C. Fan, N. Li, and Z. C. Zhou, A class of optimal ternary cyclic codes and their duals, Finite Fields Appl. 37 (2016) 193-202.

\bibitem{FL07} K. Feng and J. Luo, Value distribution of exponential sums from
perfect nonlinear functions and their applications,  IEEE Trans. Inform.
Theory 53 (9) (2007) 3035--3041.

\bibitem{TFeng} T. Feng, On cyclic codes of length $2^{2^r}-1$ with two zeros whose
dual codes have three weights, Des. Codes Cryptogr 62 (2012) 253--258.

\bibitem{Huff} W. C. Huffman and V. Pless, Fundamentals of Error-Correcting Codes, Cambridge University Press, Cambridge, 2003.

\bibitem{KL} D. J. Katz and P. Langevin, Proof of a conjectured three-valued family of Weil sums of binomials, Acta Arith. 169 (2) (2015) 181-199.

\bibitem{Klove} T. Kl{\o}ve, Codes for Error Detection, World Scientific, 2007.

\bibitem{Li-CJ}  C. J.  Li,  Q. Yue, and F. W.  Li, Weight distributions of cyclic codes with respect to pairwise coprime order elements, Finite Fields Appl. 28 (2014) 94--114.

\bibitem{Li-CL} C. L.  Li, N. Li, T. Helleseth, and C. Ding, The weight distributions of several classes of cyclic codes
from APN monomials, IEEE Trans. Inform. Theory 60 (8) (2014) 4710--4721.

\bibitem{Li-FW}  F. W.  Li, Q. Yue, and C. J.  Li, The minimum Hamming distances of irreducible cyclic codes,  Finite Fields Appl. 29 (2014) 225--242.

\bibitem{LiN} N. Li, C. Li, T. Helleseth, C. Ding, and X.H. Tang,
Optimal ternary cyclic codes with minimum distance four and five, Finite Fields Appl. 30 (2014) 100-120.

\bibitem{LZH} N. Li, Z. C. Zhou and T. Helleseth, On a conjecture about a class of optimal ternary cyclic codes, Seventh International workshop on Signal Design and its
Applications in Communications, 2015, 62-65.

\bibitem{Lint} J. H. van Lint, Introduction to Coding Theory, 3rd ed. Springer-Verlag, 1999.

\bibitem{Schmidt} B. Schmidt and C. White, All two-weight irreducible cyclic codes, Finite
Fields Appl. 8 (2002) 1--17.

\bibitem{WangLisha} L. S. Wang and G. F. Wu, Several classes of optimal ternary cyclic codes with minimal distance four, Finite Fields Appl. 40 (2016) 126-137.

\bibitem{Haode} H. D. Yan, Z. C. Zhou, and X. N. Du, A family of optimal ternary cyclic codes from the Niho-type exponent, Finite Fields Appl. 54 (2018) 101-112.

\bibitem{YangJing} J. Yang, M. Xiong, C. Ding, and J. Luo, Weight distribution of a class of cyclic codes with arbitrary number of zeros, IEEE Trans.Inform. Theory 59 (9) (2013) 5985--5993.

\bibitem{YCD} J. Yuan, C. Carlet, and C. Ding, The weight distribution of a class of linear codes from perfect nonlinear functions, IEEE Trans. Inform. Theory  52 (2) (2006) 712--717.


\bibitem{ZhengDB}D. B. Zheng, X. Q. Wang, H. Hu, and X. Zeng, The weight distributions of two classes of $p$-ary cyclic codes, Finite Fields Appl. 29 (2014) 202--242.

\bibitem{ZengXY2} X. Zeng, J. Shan,  and L. Hu, A triple-error-correcting cyclic code from the Gold and Kasami-Welch APN power functions, Finite Fields Appl. 16 (1) (2012) 70--92.

\bibitem{ZD13} Z. C. Zhou and C. Ding, A class of three-weight cyclic codes, Finite Fields Appl. 25 (2014) 79--93.

\bibitem{ZhouDingTC} Z. C. Zhou and C. Ding, Seven classes of three-weight cyclic codes, IEEE Trans. Commun.  61 (10) (2013) 4120--4126.





\end{thebibliography}

\documentclass[1p,final]{elsarticle}
\usepackage{amsfonts,natbib,pslatex}

\usepackage{latexsym,amsmath,amsthm,amssymb,amsxtra,mathrsfs,times,CJK}

\usepackage{color}
\usepackage{amsfonts,color}
\usepackage{amssymb,amsmath,latexsym}
\usepackage{amssymb}

\usepackage[para]{threeparttable}

\newcommand{\gf}{{\mathbb{GF}}}

\newcommand{\C}{{\mathcal C}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{open}[theorem]{Open Problem}

\journal{Finite Fields and Their Applications}

\begin{document}

\begin{frontmatter}

%% Title, authors and addresses

%% use the tnoteref command within \title for footnotes;
%% use the tnotetext command for the associated footnote;
%% use the fnref command within \author or \address for footnotes;
%% use the fntext command for the associated footnote;
%% use the corref command within \author for corresponding author footnotes;
%% use the cortext command for the associated footnote;
%% use the ead command for the email address,
%% and the form \ead[url] for the home page:
%%
%% \title{Title\tnoteref{label1}}
%% \tnotetext[label1]{}
%% \author{Name\corref{cor1}\fnref{label2}}
%% \ead{email address}
%% \ead[url]{home page}
%% \fntext[label2]{}
%% \cortext[cor1]{}
%% \address{Address\fnref{label3}}
%% \fntext[label3]{}

\title{On an open problem about a class of optimal ternary cyclic codes}
\author[SWJTU]{Dongchun Han}
 \ead{han-qingfeng@163.com}
\author[SWJTU]{Haode Yan\corref{cor1}}
 \ead{hdyan@swjtu.edu.cn}

 \cortext[cor1]{Corresponding author}
 \address[SWJTU]{School of Mathematics, Southwest Jiaotong University, Chengdu, 610031, China}
%\address[NNU]{College of Mathematics and Statistics, Northwest Normal University, Lanzhou, 730070, China}
\tnotetext[fn1]{D. Han's research was supported by the National Science Foundation of China Grant No.11601448 and the Fundamental Research Funds for the Central Universities of China under Grant 2682016CX121. H. Yan's research was supported by the National Natural Science Foundation of
China Grant No.11801468 and the Fundamental Research Funds for the Central Universities of China under Grant 2682018CX61.}

% use optional labels to link authors explicitly to addresses:
% \author[label1,label2]{<author name>}
% \address[label1]{<address>}
% \address[label2]{<address>}

%\author[SWJTU]{Cuiling Fan\corref{cor1}}
% \ead{cuilingfan@163.com}
%\author[UB]{Nian Li}
% \ead{nianli.2010@gmail.com}
%\author[SWJTU]{Zhengchun Zhou}
% \ead{zzc@home.swjtu.edu.cn,zczhou@126.com}
%

 %\cortext[cor1]{Corresponding author}
% \address[SWJTU]{School of Mathematics, Southwest Jiaotong University, Chengdu, 610031, China}
%\address[UB]{Department of Informatics, University of Bergen, N-5020 Bergen, Norway}
%Department of Informatics, University of Bergen, N-5020 Bergen, Norway
\begin{abstract}
Cyclic codes are a subclass of linear codes and have applications in consumer electronics, data storage systems and communication systems as they have efficient encoding and decoding algorithms. In this paper, we settle an open problem about a class of optimal ternary cyclic codes which was proposed by Ding and Helleseth \cite{Ding-Heelseth}. Let $\C_{(1,e)}$ be a cyclic code of length $3^m-1$ over $\gf(3)$ with two nonzeros $\alpha$ and $\alpha^e$, where $\alpha$ is a generator of $\gf(3^m)^*$ and $e$ is a given integer. It is shown that $\C_{(1,e)}$ is optimal with parameters $[3^m-1,3^m-1-2m,4]$ if one of the following conditions is met. 1) $m\equiv0(\mathrm{mod}~  4)$, $m\geq 4$, and $e=3^\frac{m}{2}+5$. 2) $m\equiv2(\mathrm{mod}~  4)$, $m\geq 6$, and $e=3^\frac{m+2}{2}+5$.

\end{abstract}

\begin{keyword}

Cyclic code\sep optimal code \sep ternary code\sep  Sphere Packing bound.
\MSC  94B15\sep 11T71

\end{keyword}

\end{frontmatter}

\begin{abstract}
Cyclic codes are a subclass of linear codes and have applications in consumer electronics,
data storage systems, and communication systems as they have efficient encoding and
decoding algorithms. In this paper, a conjecture proposed by Ding and Helleseth in 2013 about a class of optimal ternary cyclic codes $\C_{(1,e)}$ for $e=3^h+5, 2\leq h\leq m-1$ with parameters $[3^m-1,3^m-1-2m,4]$ is settled, if one of the following conditions is met:
\begin{enumerate}
\item $m\equiv0(\mod 4)$, $m\geq 4$, and $h=\frac{m}{2}$.
\item $m\equiv2(\mod 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.
\end{enumerate}
\end{abstract}

%\begin{keywords}
%Cyclic codes, weight distribution, quadratic form, sphere packing bound.
%\end{keywords}

\section{Introduction}

Cyclic codes are an important subclass of linear codes and have been extensively studied \cite{Klove}. Let $p$ be a prime, $m$ be a positive integer. Let $\gf(p)$ and $\gf(p^m)$ denote the finite fields with $p$ and $p^m$ elements, respectively. A linear $[n,k,d]$ code $\C$ over the finite field $\gf(p)$ is a $k$-dimensional subspace of $\gf(p)^n$ with minimum Hamming distance $d$, and is called {cyclic} if any cyclic shift of a codeword
is another codeword of $\C$. Let $\gcd(n,p)=1.$ By identifying any vector $(c_0,c_1,\cdots,c_{n-1})\in \gf(p)^n$ with
 \[c_0+c_1x+c_2x^2+\cdots+c_{n-1}x^{n-1}\in\gf(p)[x]/(x^n-1),\]
 any cyclic code of length $n$ over $\gf(p)$ corresponds to an ideal of the polynomial residue class ring $\gf(p)[x]/(x^n-1)$.
 It is well known that every ideal of $\gf(p)[x]/(x^n-1)$ is principal. Any cyclic code $\C$ can be expressed as $\C=\langle g(x) \rangle$, where $g(x)$ is monic and has the least degree. Then $g(x)$ is called the {generator polynomial} and $h(x)=(x^n-1)/g(x)$ is referred to as the { parity-check polynomial} of $\C$. For some recent developments of cyclic codes, the readers are referred to \cite{CDY05}, \cite{Ding12}-\cite{Ding-Heelseth}, \cite{FLZ}, \cite{TFeng}, \cite{Li-CJ}-\cite{LiN},  \cite{Schmidt}-\cite{YangJing}, \cite{ZhengDB}-\cite{ZhouDingTC}  and the references therein.

Let $\alpha$ be a generator of $\gf(3^m)^*=\gf(3^m)\setminus\{0\}$ and $m_{i}(x)$ be
the minimal polynomial of $\alpha^i$ over $\gf(3)$, where $1\leq i\leq 3^m-1$.
Let $\C_{(1,e)}$ be the cyclic code over $\gf(3)$  with generator polynomial $m_1(x)m_e(x)$, where $e$ is an integer such that
$\alpha$ and $\alpha^e$ are nonconjugate. Carlet, Ding and Yuan \cite{CDY05} proved that $\C_{(1,e)}$ has parameters $[3^m-1,3^m-1-2m,4]$ when $x^e$ are certain perfect nonlinear monomials over $\gf(3^m)$. Notice that the ternary cyclic code with parameters $[3^m-1,3^m-1-2m,4]$ is optimal according to the Sphere Packing bound. In 2013, Ding and Helleseth \cite{Ding-Heelseth} constructed several classes of
optimal ternary cyclic codes $\C_{(1,e)}$ with parameters $[3^m-1,3^m-1-2m,4]$ by employing some monomials $x^e$ over $\gf(3^m)$ including almost perfect nonlinear monomials. In addition, nine open problems about $\C_{(1,e)}$ with parameters $[3^m-1,3^m-1-2m,4]$ were proposed in \cite{Ding-Heelseth}. Recently, two of the nine open problems were solved, see \cite{LiN, LZH}. Moreover, an open problem proposed in \cite{Ding-Heelseth} is shown as follows.

\begin{open}[Open Problem 7.12, \cite{Ding-Heelseth}]\label{open-1}
Let $e=3^h+5$, where $2\leq h\leq m-1$. Let $m$ be even. Is it true that the ternary cyclic code $\C_{(1,e)}$ has parameters  $[3^m-1,3^m-1-2m,4]$ if one of the following conditions is met?
\begin{enumerate}
\item $m\equiv0(\mathrm{mod}~  4)$, $m\geq 4$, and $h=\frac{m}{2}$.

\item $m\equiv2(\mathrm{mod}~  4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.

\end{enumerate}
\end{open}

In this paper, we will settle this open problem. The rest of this paper is organized as follows. In section \ref{preliminaries}, we introduce two useful results which will
be employed in the sequel. In Section \ref{main}, we present the proof of our main result. Section \ref{conclusion} concludes the paper with some remarks.

\section{Preliminaries}\label{preliminaries}

In this section, we will introduce two useful results. The first one is about the cyclotomic coset. For a prime $p$, the $p$-cyclotomic coset modulo $p^m-1$ containing $j$ is defined as
$$\C_j=\big\{jp^s \mod(p^m-1):s=0,1,...,m-1\big\}.$$
We have the following lemma.
\begin{lemma}[Lemma 2.1, \cite{Ding-Heelseth}]\label{Lemma-size}
For any $1\leq e\leq p^m-2$ with $\gcd(e,p^m-1)=2$, the cardinality of the $p$-cyclotomic coset $\C_e$ is equal to $m$.
\end{lemma}

%\begin{lemma}\label{Lemma-irr}
%For every finite field $\gf(q)$ and every positive integer $n$, where $q$ is a prime power, the product of all monic irreducible polynomials over $\gf(q)[x]$ whose degrees divide $n$ is equal to $x^{q^n}-x$.
%\end{lemma}
It is known that a code with parameters $[3^m-1,3^m-1-2m,4]$ is optimal according to the Sphere Packing bound. To determine the optimality of $\C_{(1,e)}$, the following sufficient and necessary conditions are
given by Ding and Helleseth in \cite{Ding-Heelseth}.

\begin{theorem}[Theorem 4.1, \cite{Ding-Heelseth}]\label{thm-fundamental}
Let $e\notin\C_1$, and $|\C_e|=m$. The ternary cyclic code $\C_{(1,e)}$ has parameters  $[3^m-1,3^m-1-2m,4]$ if and only if the following conditions are satisfied:

{\textit{C1}}: e is even;

{\textit{C2}}: the equation $(x+1)^e-x^e-1=0$ has the only solution $x=0$ in $\mathsf \gf(3^m)$;

{\textit{C3}}: the equation $(x+1)^e+x^e+1=0$ has the only solution $x=1$ in $\mathsf \gf(3^m)$.

\end{theorem}

\section{Solving Open Problem \ref{open-1}}\label{main}
In this section, we confirm that each condition in Open Problem \ref{open-1} satisfies all the three conditions in Theorem \ref{thm-fundamental}. Then the answer of the open problem can be deduced. Firstly, we confirm that C1 holds in the following lemma.
\begin{lemma}\label{Lemma-size1}
Let $e=3^h+5$, where $2\leq h\leq m-1$. Then $e\notin\C_1$ and $|\C_e|=m$ if one of the following conditions is met.
\begin{enumerate}
\item $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

\item $m\equiv2(\mathrm{mod}~ 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.

\end{enumerate}
\end{lemma}
\begin{proof}
We only prove the first one and the second one is similar. It is easy to see that $e\notin\C_1$ since $e$ is even. It will be shown that $|\C_e|=m$.
We have
\begin{align*}
\gcd(e, 3^m-1)&=\gcd(3^{\frac{m}{2}}+5,3^m-1)=\gcd(3^{\frac{m}{2}}+5,3^m-1-(3^{\frac{m}{2}}+5)(3^{\frac{m}{2}}-5))\\
     &=\gcd(3^{\frac{m}{2}}+5,24)=\gcd(3^{\frac{m}{2}}+5,8)=(6,8)=2.
\end{align*}
The fifth equality holds since $m\equiv0(\mathrm{mod}~ 4)$ and $3^{\frac{m}{2}}+5 \equiv 6 (\mathrm{mod}~ 8)$. Consequently, $|\C_e|=m$ follows from lemma \ref{Lemma-size}.
\end{proof}
Secondly, we investigate the solutions of $(x+1)^e-x^e-1=0$ in $\gf(3^m)$.
\begin{lemma}\label{Lemma-C2}
Let $e=3^h+5$, where $2\leq h\leq m-1$. Then
\begin{equation}\label{eqnminus}(x+1)^e-x^e-1=0
\end{equation}
has the only solution $x=0$ in $\mathsf \gf(3^m)$ if one of the following conditions is met.
\begin{enumerate}
\item $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

\item $m\equiv2(\mathrm{mod}~ 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.
\end{enumerate}
\end{lemma}
\begin{proof}
It is obvious that $x=0$ is a solution of (\ref{eqnminus}) and $x=\pm 1$ is not. Suppose that $\theta\in \gf(3^m)\setminus \gf(3)$ and is a solution of (\ref{eqnminus}). Through a straight calculation, we have that
$$\theta^{3^h-1}(\theta^4-\theta^3-\theta^2+\theta-1)=\theta^4-\theta^3+\theta^2+\theta-1.$$
First, we assert that $\theta^4-\theta^3-\theta^2+\theta-1\neq0$. Otherwise, we have $\theta^4-\theta^3-\theta^2+\theta-1=\theta^4-\theta^3+\theta^2+\theta-1=0$, which leads to $\theta=0$. It is a contradiction. Hence we have
\begin{eqnarray}\label{eqn-solution1}
\theta^{3^{h}}=\frac{f(\theta)}{g(\theta)},
\end{eqnarray}
where $f(\theta)=\theta^5-\theta^4+\theta^3+\theta^2-\theta$ and $g(\theta)=\theta^4-\theta^3-\theta^2+\theta-1$.
Taking $3^{h}$ powers on both sides of the equation (\ref{eqn-solution1}), we have
\begin{equation}\label{eqn-solution2}
\theta^{3^{2h}}=\frac{\theta^{5\cdot3^h}-\theta^{4\cdot3^h}+\theta^{3\cdot3^h}+\theta^{2\cdot3^h}-\theta^{3^h}}{\theta^{4\cdot3^h}-\theta^{3\cdot3^h}-\theta^{2\cdot3^h}+\theta^{\cdot3^h}-1}.
\end{equation}
Plugging (\ref{eqn-solution1}) into (\ref{eqn-solution2}), we obtain
\begin{equation}\label{eqn-solution2.5}
\theta^{3^{2h}}=\frac{F(\theta)}{G(\theta)},
\end{equation}
where
$F(\theta)=f(\theta)^5-f(\theta)^4g(\theta)+f(\theta)^3g(\theta)^2+f(\theta)^2g(\theta)^3-f(\theta)g(\theta)^4$ and
$G(\theta)=f(\theta)^4g(\theta)-f(\theta)^3g(\theta)^2-f(\theta)^2g(\theta)^3+f(\theta)g(\theta)^4-g(\theta)^5$.
We distinguish the following two cases.

{\textit{Case 1}}: $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

Noting that $\theta^{3^{2h}}=\theta$ since $2h=m$, then (\ref{eqn-solution2.5}) becomes
$$\theta G(\theta) - F(\theta)=0.$$
With the help of Magam Program, we can decompose the left-hand side of the above equation into the product of some irreducible factors as follows.
$$\theta^3(\theta+1)(\theta-1)(\theta^6+\theta^3-\theta+1)(\theta^6-\theta^5+\theta^3+1)(\theta^6-\theta^5-\theta^3-\theta+1)=0.$$

If $\theta^6+\theta^3-\theta+1=0$, then $\theta \in \gf(3^6)\subseteq\gf(3^m)$. We have $6|m$ and then $6|h$ since $m=2h$ and $h$ is even. Plugging $\theta^{3^h}=\theta$ into the equation (\ref{eqn-solution1}), we have
$$\theta=\frac{\theta^5-\theta^4+\theta^3+\theta^2-\theta}{\theta^4-\theta^3-\theta^2+\theta-1},$$
which leads to $\theta=0$. It is a contradiction. Similarly, we can prove that $\theta^6-\theta^5+\theta^3+1\neq0$ and  $\theta^6-\theta^5-\theta^3-\theta+1\neq0$. Then $x=0$ is the only solution of (\ref{eqnminus}) in $\gf(3^m)$.

{\textit{Case 2}}: $m\equiv2(\mathrm{mod}~  4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.

Noting that $\theta^{3^{2h}}=\theta^9$ since $2h=m+2$, then (\ref{eqn-solution2.5}) becomes
$$\theta^9 G(\theta) - F(\theta)=0.$$

With the help of Magam Program, we can decompose the left-hand side of the above equation into the product of some irreducible factors as follows.
$$\theta(\theta+1)(\theta-1)(\theta^4+\theta^3-\theta^2-\theta-1)(\theta^4+\theta^3+\theta^2-\theta-1)(\theta^6-\theta^5+\theta^4-\theta^3+\theta^2-\theta+1)$$
$$(\theta^8+\theta^7+\theta^6-\theta^4+\theta^2+\theta+1)(\theta^8+\theta^7-\theta^6-\theta^2+\theta+1)=0.$$

If  $\theta^6-\theta^5+\theta^4-\theta^3+\theta^2-\theta+1=0$, then $\theta \in \gf(3^6)\subseteq\gf(3^m)$ and $6|m$. It follows from $h$ is even that $h\equiv4(\mathrm{mod}~ 6)$. Noting that $\theta^7=-1$, we obtain $\theta^{3^h}=\theta^{3^4}=\theta^{77+4}=-\theta^4$. Plugging this into (\ref{eqn-solution1}), we obtain
$$\theta^6+\theta^5+\theta^4-\theta^3-\theta^2-\theta-1=0.$$
This together with $\theta^6-\theta^5+\theta^4-\theta^3+\theta^2-\theta+1=0$ leads to $\theta^5-\theta^2-1=\theta^5+\theta^3-\theta^2-1=0$. Then $\theta$ must be zero, which is a contradiction. Moreover,
$\theta^4+\theta^3-\theta^2-\theta-1, \theta^4+\theta^3+\theta^2-\theta-1, \theta^8+\theta^7+\theta^6-\theta^4+\theta^2+\theta+1$ and $\theta^8+\theta^7-\theta^6-\theta^2+\theta+1$ cannot be zero since $4\nmid m$. Then $x=0$ is the only solution of (\ref{eqnminus}) in $\gf(3^m)$.
This completes the proof.
\end{proof}
In what follows, we investigate the solutions of $(x+1)^e+x^e+1=0$ in $\gf(3^m)$.
\begin{lemma}\label{Lemma-C3}
Let $e=3^h+5$, where $2\leq h\leq m-1$. Then
\begin{equation}\label{eqnplus}
(x+1)^e+x^e+1=0
\end{equation} has the only solution $x=1$ in $\mathsf \gf(3^m)$ if one of the following conditions is met.
\begin{enumerate}
\item $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

\item $m\equiv2(\mathrm{mod}~ 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.
\end{enumerate}
\end{lemma}
\begin{proof}
It is obvious that $x=1$ is a solution of (\ref{eqnplus}). Suppose that $\theta\in\gf(3^m)\setminus \gf(3)$ is a solution of (\ref{eqnplus}). Through a straight calculation, we have
$$\theta^{3^h}(\theta^4-\theta^3+\theta^2+1)=\theta^4+\theta^2-\theta+1.$$
First, we assert that $\theta^4-\theta^3+\theta^2+1\neq0$. Otherwise, we have $\theta^4-\theta^3+\theta^2+1=\theta^4+\theta^2-\theta+1=0$. It then follows that $\theta^3=\theta$. This is contrary to the assumption that $\theta \notin \gf(3)$. Hence, we have
\begin{eqnarray}\label{eqn-solution3}
\theta^{3^{h}}=\frac{k(\theta)}{l(\theta)},
\end{eqnarray}
where $k(\theta)=\theta^4+\theta^2-\theta+1$ and $l(\theta)=\theta^4-\theta^3+\theta^2+1$. Taking $3^{h}$ powers on both sides of the equation (\ref{eqn-solution3}), we have
\begin{equation}\label{eqn-solution4}
\theta^{3^{2h}}=\frac{\theta^{4\cdot3^h}+\theta^{2\cdot3^h}-\theta^{3^h}+1}{\theta^{4\cdot3^h}-\theta^{3\cdot3^h}+\theta^{2\cdot3^h}+1},
\end{equation}
Plugging (\ref{eqn-solution3}) into (\ref{eqn-solution4}), we obtain
\[\theta^{3^{2h}}=\frac{K(\theta)}{L(\theta)},\]
where $K(\theta)=k(\theta)^4+k(\theta)^2l(\theta)^2-k(\theta)l(\theta)^3+l(\theta)^4$ and $L(\theta)=k(\theta)^4-k(\theta)^3l(\theta)+k(\theta)^2l(\theta)^2+l(\theta)^4$. We distinguish the following two cases.

{\textit{Case 1}}: $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

Noting that $\theta^{3^{2h}}=\theta$ since $2h=m$, then $\theta$ satisfies
$$\theta L(\theta)-K(\theta)=0.$$
With the help of Magam Program, we can decompose the left-hand side of the above equation into the product of some irreducible factors as follows.
$$(\theta-1)^5(\theta^2+\theta-1)^2(\theta^2-\theta-1)^2(\theta^2+1)^2=0.$$

If $\theta^2+\theta-1=0$, then $\theta \in \gf(3^2)$. We have $\theta^{3^2}=\theta$ and then $\theta^{3^h}=\theta$ since $h$ is even. Plugging $\theta^{3^h}=\theta$ into (\ref{eqn-solution3}), then we have
$$\theta=\frac{\theta^4+\theta^2-\theta+1}{\theta^4-\theta^3+\theta^2+1},$$
which leads to $\theta^5+\theta^4+\theta^3-\theta^2-\theta-1=(\theta-1)^5=0$, a contradiction. Similarly, we can prove that $\theta^2+\theta-1\neq0$ and $\theta^2+1\neq0$. Then $x=1$ is the only solution of (\ref{eqnplus}) in $\gf(3^m)$.

{\textit{Case 2}}: $m\equiv2(\mathrm{mod}~ 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.

Noting that $\theta^{3^{2h}}=\theta^9$ since $2h=m+2$, then $\theta$ satisfies
$$\theta^9 L(\theta)-K(\theta)=0.$$
With the help of Magam Program, we can decompose the left-hand side of the above equation into the product of some irreducible factors as follows.
$$(\theta-1)(\theta^2+1)(\theta^2+\theta-1)(\theta^2-\theta-1)(\theta^3-\theta+1)(\theta^3-\theta-1)$$
\[(\theta^3+\theta^2-\theta+1)(\theta^3-\theta^2+\theta+1)(\theta^3+\theta^2-1)(\theta^3-\theta^2+1)=0.\]
Similar with the proof of Case 1, we know that $\theta^2+1,\theta^2+\theta-1,\theta^2-\theta-1\neq0$. If $\theta^3-\theta-1=0$, then $\theta \in \gf(3^3)\subseteq \gf(3^m)$. We have $\theta^{3^3}=\theta$ and $3|m=2h-2$, this leads to $h\equiv1(\mathrm{mod}~ 3)$ and $\theta^{3^h}=\theta^3$. Plugging $\theta^{3^h}=\theta^3$ into (\ref{eqn-solution3}), we have
$$\theta^3=\frac{\theta^4+\theta^2-\theta+1}{\theta^4-\theta^3+\theta^2+1},$$
which leads to $\theta^8=1$. It follows from $\theta^{3^3-1}=1$ that $\theta^2=\theta^{(8,3^3-1)}=1$, a contradiction. This completes the proof of Case $2$.
\end{proof}

The answer to Open Problem \ref{open-1} is given in the following theorem.

\begin{theorem}
Let $e=3^h+5$, where $2\leq h\leq m-1$. Let $m$ be even. Then the ternary cyclic code $\C_{(1,e)}$ has parameters  $[3^m-1,3^m-1-2m,4]$ if one of the following conditions is met.
\begin{enumerate}
\item $m\equiv0(\mathrm{mod}~ 4)$, $m\geq 4$, and $h=\frac{m}{2}$.

\item $m\equiv2(\mathrm{mod}~ 4)$, $m\geq 6$, and $h=\frac{m+2}{2}$.

\end{enumerate}
\end{theorem}
\begin{proof}
The conclusions follow from Lemma \ref{Lemma-size1}, Lemma \ref{Lemma-C2}, Lemma \ref{Lemma-C3} and Theorem \ref{thm-fundamental}.
\end{proof}

\section{Conclusions}\label{conclusion}

In this paper, we settled an open problem proposed by Ding and Helleseth in 2013 about a class of optimal ternary cyclic codes. The main technique we used is shown in solving the equation in conditions C2 and C3. Assume that $\theta$ is a solution of the target equation, we can obtain $\theta^{3^h}=R(\theta)$ after calculation, where $R(\theta)$ is a rational function of $\theta$ with known degree and coefficients. Then we take $3^h$-th power of $\theta^{3^h}=R(\theta)$, together with the relationship between $m$ and $h$, we can find an solvable equation of $\theta$. We remark that when $h$ is close to $\gamma m$, where $\gamma$ is a rational number, our technique always works. For instance, the following theorem gives other optimal cyclic codes with respect to the Sphere Packing bound. This gives an incomplete answer to Open Problems 7.12-7.15 in \cite{Ding-Heelseth}.
\begin{theorem}
Let $m$ be an odd integer no less than five and $\gcd(m,3)=1$. Then the ternary cyclic code $\C_{(1,e)}$ has parameters  $[3^m-1,3^m-1-2m,4]$ if one of the following conditions is met.
\begin{enumerate}
\item $e=3^h+5$, where $2h\equiv\pm1(\mathrm{mod}~ m)$;

\item $e=3^h+13$, where $2h\equiv\pm1(\mathrm{mod}~ m)$;

\item $e=\frac{3^{m-1}}{2}+3^h+1$, where $2h\equiv\pm1(\mathrm{mod}~ m)$ or $3h\equiv\pm1(\mathrm{mod}~ m)$ or $4h\equiv\pm1(\mathrm{mod}~ m)$.
\end{enumerate}
\end{theorem}

%\section*{Acknowledgments}
%The authors are very grateful to the reviewer and the
%Associate Editor, Prof. James W.P. Hirschfeld, for their comments and
%suggestions that improved the presentation and quality of this
%paper. This work was finished when the authors visited the Hong Kong University of Science and Technology. The authors
%are grateful to Professor Cunsheng Ding for bringing them together
%in the summer of 2014.

%\section*{Acknowledgements}
%The authors are grateful to the anonymous reviewers for careful reading and for invaluable suggestions which improve the quality of the paper.

\end{document}